a certain machine makes electrical resistors having a mean resistance of 40 ohms and standard deviation of 2 ohms what is the probability that random sample of 36 have combined resistance more than 1458


1
Expert's answer
2022-06-13T15:13:51-0400
"1458\/36=40.5"


"P(X>40.5)=1-P(Z\le\dfrac{40.5-40}{2/\sqrt{36}})"

"=1-P(Z\le1.5)\approx0.0668"